2,24,46 ... Find the 25th term
42,32,22,12 ... Find the 35th term
3,-18,108,-648 ... Find the 22nd term
2,1,1/2,1/4 ... Find the 9th term
Someone please help me with these? I need this worksheet done asap

Answer by stanbon(75887)

2,24,46 ... Find the 25th term
a(1) = 2 ; common difference = d = 22
a(25) = a(1)+(25-1)d = 2+24*22 = 530
----------------------------------------
42,32,22,12 ... Find the 35th term
a(1) = 42 ; d = -10
a(35) = 42 + 34(-10) = -298
-------
3,-18,108,-648 ... Find the 22nd term
a(1) = 3 ; common ratio = -18/3 = r = -6
a(22) = a(1)*r^(n-1) = 3*(-6)^21 = -6.58108519*10^16
-----------------------------
2,1,1/2,1/4 ... Find the 9th term
a(1) = 2 ; r = 1/2
a(9) = 2*(1/2)^8 = 2^7 = 128
-------------
Cheers,
Stan H.
--------------